# RRWNet: Recursive Refinement Network for Effective Retinal Artery/Vein Segmentation and Classification

### v3

@@ 6 Results and Discussion
6.1 Hyperparameter search
@@
Table 2: Impact of $K$ in RITE (validation). Proposed RRWNet with different $K$: 2, 3, 6, 8, and 11. Best results are highlighted in bold, and second best results are underlined. All values are percentages.
@@
Fig. 6 displays the segmentation maps generated by RRWNet ($K=6$) at each iteration $k$.
@@
As evident in the figure, the RR module progressively improves A/V classification over iterations. Notably, in the base segmentation map ($k=0$), several misclassified vessels are visible. However, these errors are progressively reduced in subsequent iterations, resulting in a final segmentation map ($k=6$) with significantly fewer manifest classification errors and improved delineation of both arteries and veins.
@@ 6.2 Ablation study
Table 3: Ablation study in RITE. W-Net: 2 stacked U-Nets without recursive refinement ($K=1$), similar to Galdran et al., [2022]. RRU-Net: recursive refinement using a single U-Net module, similar to Mosinska et al., [2018]. RRWNetAll: proposed approach with recursive refinement ($K=6$) for arteries, veins, and BV. RRWNet: proposed approach with recursive refinement of arteries and veins only ($K=6$). A one-tailed Wilcoxon signed-rank test was performed to compare the results of the proposed RRWNet with the results of the best or second best model for each evaluation metric. *: $p<0.05$. Best results are highlighted in bold, and second best results are underlined.
@@
Similar to segmentation, all methods surpassed the U-Net baseline in all A/V classification metrics, with RRWNet demonstrating statistically significant superiority over the second-best method in all cases. Notably, RRWNet outperformed U-Net by significant margins in sensitivity (+7.41 pp), specificity (+3.89 pp), and accuracy (+5.48 pp). Similar to RRWNet, RRWNetAll achieved improved A/V classification performance compared to W-Net and RRU-Net, exhibiting an accuracy increase of 1.83 pp and 2.35 pp, respectively. Smaller improvements in A/V accuracy were observed for both W-Net (+1.54 pp) and RRU-Net (+1.02 pp) compared to the U-Net baseline.
@@
6.3 State of the art comparison
@@ 6.3.2 A/V segmentation and classification for all GT vessels and RR post-processing
In addition to the standard state-of-the-art comparison presented in Table 4, Table 5 offers a more comprehensive analysis of the performance of the proposed model compared to existing approaches. In particular, for A/V classification, the comparison is performed in terms of accuracy (Acc.), considering all “vessel” pixels from the GT except crossings and unknown pixels For A/V segmentation, the comparison is performed using different threshold-agnostic metrics (AUPR and AUROC) and topological metrics (COR and INF). The table also includes, in parentheses, the results obtained by applying the proposed RR module as a post-processing step to the segmentation maps generated by the other methods. This provides insight into the potential benefit of the RR module for enhancing existing approaches.
Table 5: Comparison with the state of the art in artery and vein classification and segmentation. In this case, all “vessel” pixels from the GT except crossings and unknown pixels are considered for the evaluation. All values are in percentages. The results obtained by applying the proposed RR module as a post-processing step to the segmentation maps generated by the other methods are shown in parentheses. When this value is higher than the value obtained by the method itself, it is highlighted in green; when it is lower, in red. The best values among the automatic end-to-end methods (i.e., excluding the post-processing step) for each metric and dataset are highlighted in bold. The best overall (including both with and without the post-processing step) are underlined.
@@
RRWNet (ours) 95.85 84.99 98.91 48.40 51.00 88.36 98.99 48.13 51.40
RRWNet consistently outperformed state-of-the-art methods on all datasets and metrics considered in Table 5. In RITE, RRWNet achieves 94.95% A/V accuracy, outperforming all other methods by at least 0.28 pp. Similar improvements are observed in AUPR and AUROC for both artery and vein segmentation. However, the most significant improvements are observed for the metrics measuring topological consistency: COR and INF. RRWNet achieves 31.62% COR and 68.03% INF for arteries and 38.23% COR and 61.36% INF for veins. These values represent substantial advancements, with COR being 12.58 pp higher and INF 12.53 pp lower (for INF, lower is better) for arteries and 11.69 pp higher and 10.87 pp lower for veins compared to the second-best method. This underlines the superior ability of RRWNet to generate more topologically correct segmentation maps compared to the state of the art. The results are similar for the HRF dataset, and even more remarkable for the LES-AV dataset, where RRWNet greatly outperforms the state-of-the-art methods in all metrics. The differences are, again, particularly pronounced for the topological metrics, with RRWNet achieving 47.05% COR and 51.68% INF for arteries (+35.31 pp and -36.14 pp over the second-best method, respectively) and 49.68% COR and 49.45% INF for veins (+23.77 pp and -24.46 pp over the second-best method, respectively). This showcases the generalization capabilities of RRWNet to different an unseen datasets, and its salient ability to generate topologically correct segmentation maps.
@@
Examples of segmentation maps obtained by the different models compared in Table 5 are shown in Figs. 8 and 9, for RITE and HRF datasets, respectively [^10] All the segmentation maps obtained by our RRWNet model for RITE, HRF and LES-AV datasets will be publicly available at https://github.com/j-morano/rrwnet..
@@ 7 Conclusions
To rigorously assess the efficacy of the proposed framework, we implemented a straightforward instantiation based on the well-established U-Net architecture. This implementation was evaluated on the task of A/V segmentation and classification within several publicly available retinography image datasets. The quantitative results demonstrated that the proposed method outperformed state-of-the-art methods by a notable margin, both in terms of A/V classification accuracy and, more remarkably, of topological consistency. Furthermore, the standalone application of the proposed RR module demonstrably improved the segmentation maps generated by all the other compared methods, further substantiating its effectiveness as a generalizable post-processing step.
@@
In conclusion, the proposed framework and its implementation represent an effective approach to A/V segmentation and classification, with the potential to be extended to other semantic segmentation tasks and modalities. We believe that this work will serve as a good reference implementation and benchmark and encourage further research in this direction, and that it will contribute to the development of more robust and accurate semantic segmentation systems, with a particular focus on the field of ophthalmology.
@@ Acknowledgements
This work was supported in part by the Christian Doppler Research Association, Austrian Federal Ministry for Digital and Economic Affairs, the National Foundation for Research, Technology and Development.

### v4

@@ 6 Results and Discussion
6.1 Hyperparameter search: Number of refinement steps
@@
Table 2: Impact of $K$ in RITE (joint validation set). Proposed RRWNet with different $K$: 2, 3, 6, 8, and 11. All “vessel” pixels from the GT except crossings and unknown pixels are considered for the evaluation. All values are percentages. Best results are highlighted in bold, and second best results are underlined.
@@
6.2 Effect of the refinement module
Table 3 shows the evolution of the A/V and BV/BG classification performance of the selected RRWNet ($K=6$) over the refinement iterations on the RITE dataset.
Table 3: Evolution of the A/V and BV/BG classification performance of RRWNet (best model) in RITE over the refinement iterations. Since BV maps are not refined, BV/BG classification performance remains constant across iterations. In this case, all “vessel” pixels from the GT except crossings and unknown pixels are considered for the evaluation. All values are in percentages. The best values for each metric are highlighted in bold. The $p$-values of the Wilcoxon signed-rank test comparing the performance at $k=0$ and $k=6$ are also shown: **: $p<0.001$.
Iteration ($k$) A/V BV/BG
Acc. Sens. Spec. Acc. Sens. Spec. AUROC
0 89.62 86.39 92.26 96.29 80.16 98.61 98.50
1 93.42 92.22 94.41 = = = =
2 94.46 93.76 95.03 = = = =
3 94.78 94.19 95.26 = = = =
4 94.94 94.45 95.34 = = = =
5 94.92 94.42 95.34 = = = =
6 94.95** 94.46** 95.38** = = = =
The results show a progressive and, in the end, significant improvement in A/V classification performance over the refinement iterations. For instance, the accuracy increased significantly from 89.62% at $k=0$ to 94.95% (+5.33 pp) at $k=6$, with a corresponding increase in sensitivity and specificity. The improvement can also be observed at a glance in the segmentation maps. Fig. 6 displays the segmentation maps generated by the same model at each iteration $k$ for a sample image from the RITE dataset.
@@ 6.2 Effect of the refinement module
As evident in the figure, the RR module progressively improves the A/V classification over the iterations. Notably, in the base segmentation map ($k=0$), several misclassified vessels are visible. However, these errors are progressively reduced in subsequent iterations, resulting in a final segmentation map ($k=6$) with significantly fewer manifest classification errors and improved delineation of both arteries and veins.
@@ 6.3 Ablation study
Table 4: Ablation study in RITE. W-Net: 2 stacked U-Nets without recursive refinement ($K=1$), similar to Galdran et al., [2022]. RRU-Net: recursive refinement using a single U-Net module, similar to Mosinska et al., [2018]. RRWNetAll: proposed approach with recursive refinement ($K=6$) for arteries, veins, and BV. RRWNet: proposed approach with recursive refinement of arteries and veins only ($K=6$). The results are the average of the 4 models trained for each fold. All “vessel” pixels from the GT except crossings and unknown pixels are considered for the evaluation. A one-tailed Wilcoxon signed-rank test was performed to compare the results of the proposed RRWNet with the results of the best or second best model for each evaluation metric. *: $p<0.05$. Best results are highlighted in bold, and second best results are underlined.
@@
Similar to segmentation, all methods surpassed the U-Net baseline in all A/V classification metrics, with RRWNet demonstrating statistically significant superiority over the second-best method in all cases. Notably, RRWNet outperformed U-Net by significant margins in sensitivity (+7.41 pp), specificity (+3.89 pp), and accuracy (+5.48 pp). Similar to RRWNet, but with significantly lower performance, RRWNetAll achieved improved A/V classification performance compared to W-Net and RRU-Net, exhibiting an accuracy increase of 1.83 pp and 2.35 pp, respectively. Smaller improvements in A/V accuracy were observed for both W-Net (+1.54 pp) and RRU-Net (+1.02 pp) compared to the U-Net baseline.
The significant differences between RRWNet and RRWNetAll can be explained by the greater specialization of the RR module of the former. In RRWNet, the RR module is focused only on refining the A/V segmentation maps (by solving manifest classification errors), and not the BV maps, so its full capacity can be dedicated to this matter. Conversely, in RRWNetAll, the RR module also aims at correcting the entire vessel segmentation. However, we have observed empirically that refining all maps (and thus refining the entire segmentation, not just the A/V classification) often leads to hallucinated vessels or vessel connections (see Fig. 7). These hallucinations affect the topology of the segmentation maps, introducing noise and thus complicating the resolution of A/V manifest classification errors based on topological information. This observation is consistent with the reduced performance of the RRWNetAll model in the segmentation of BV (arteries+veins) compared to the other models (including a simple U-Net), as shown in Table 4 and illustrated in Figs. 7 and 8. All of this strongly suggests that the specialization of the RR module is a key factor in its effectiveness.
GT RRWNetAll RRWNet
Refer to caption Refer to caption Refer to caption
Figure 7: Examples of segmentation maps obtained by RRWNetAll and RRWNet. RRWNetAll tends to hallucinate blood vessels, altering the topology. Some differences between the segmentation maps are highlighted in colored boxes. [RITE, image 05]
@@
6.4 Efficiency analysis
We compare RRWNet ($k=6$), W-Net (which is equivalent to RRWNet with $k=1$), and U-Net (no refinement, $k=0$) in terms of computational efficiency on the RITE-test dataset. In particular, we evaluate the inference time per sample, number of multiply-accumulate operations (MAC), number of parameters, and memory consumption during inference, with batch size set to 1. For context, we also show the A/V classification accuracy (from Table 4) for each model. The code was run on a machine with an AMD Ryzen 7 5800X (16) @ 4.85 GHz CPU, 32 GB RAM, and an NVIDIA GeForce RTX 3060 Ti Lite Hash Rate GPU with 8 GB of memory. The results are summarized in Table 5.
Table 5: Efficiency analysis of RRWNet, W-Net, and U-Net on the RITE-test dataset. When applicable, the average and standard deviation are reported.
Method
Refinement
steps ($K$)
Inference time
on GPU (ms)
Inference time
on CPU (s)
MACs
Parameters
Memory
A/V Acc.
U-Net 0 $1.60\pm 0.07$ $1.03\pm 0.03$ 1.14 62.06 1.05 $89.14\pm 0.67$
W-Net 1 $3.01\pm 0.06$ $2.17\pm 0.07$ 2.27 124.13 1.42 $91.68\pm 0.75$
RRWNet 6 $9.25\pm 0.14$ $7.66\pm 3.49$ 7.96 124.13 1.44 $94.63\pm 0.24$
The results show that the time and MACs required for inference increase linearly with the number of refinement steps. The number of parameters for W-Net and RRWNet is the same, and twice that of U-Net, as both models contain a refinement module. However, the memory consumption of these models, which is roughly the same, is only 40% higher than that of U-Net. Despite the increased computational cost, classification accuracy improves significantly with the proposed RRWNet model, and the requirements remain acceptable for most practical applications, with an average inference time of less than 10ms on GPU and 10s on CPU, and less than 1.5GB of memory.
6.5 State-of-the-art comparison
@@ 6.5.2 A/V segmentation and classification for all GT vessels and RR post-processing
In addition to the standard state-of-the-art comparison presented in Table 6, Table 7 offers a more comprehensive analysis of the performance of the proposed model compared to existing approaches. In particular, for A/V classification, the comparison is performed in terms of accuracy (Acc.), considering all “vessel” pixels from the GT except crossings and unknown pixels. For A/V segmentation, the comparison is performed using different threshold-agnostic metrics (AUPR and AUROC) and topological metrics (COR and INF). The table also includes, in parentheses, the results obtained by applying the proposed RR module as a post-processing step to the segmentation maps generated by the other methods. This provides insight into the potential benefit of the RR module for enhancing existing approaches.
Table 7: Comparison with the state of the art in artery and vein classification and segmentation. In this case, all “vessel” pixels from the GT except crossings and unknown pixels are considered for the evaluation. In all cases, the number of refinement steps is $K=6$. All values are in percentages. The results obtained by applying the proposed RR module as a post-processing step to the segmentation maps generated by the other methods are shown in parentheses. When this value is higher than the value obtained by the method itself, it is highlighted in green; when it is lower, in red. The best values among the automatic end-to-end methods (i.e., excluding the post-processing step) for each metric and dataset are highlighted in bold. The best overall (including both with and without the post-processing step) are underlined. The $p$-values of the Wilcoxon signed-rank test (using bootstrapping) between the best and second best models are indicated as follows: *: $p<0.05$, **: $p<0.001$.
@@
RRWNet (ours) 95.85* 84.99** 98.91** 48.40** 51.00** 88.36** 98.99** 48.13** 51.40**
RRWNet consistently outperformed state-of-the-art methods on all datasets and metrics considered in Table 7 by significant differences. In RITE, RRWNet achieves 94.95% A/V accuracy, outperforming all other methods by at least 0.28 pp. Similar improvements are observed in AUPR and AUROC for both artery and vein segmentation. However, the most significant improvements are observed for the metrics measuring topological consistency: COR and INF. RRWNet achieves 31.62% COR and 68.03% INF for arteries and 38.23% COR and 61.36% INF for veins. These values represent substantial advancements, with COR being 12.58 pp higher and INF 12.53 pp lower (for INF, lower is better) for arteries and 11.69 pp higher and 10.87 pp lower for veins compared to the second-best method. This underlines the superior ability of RRWNet to generate more topologically correct segmentation maps compared to the state of the art. The results are similar for the HRF dataset, and even more remarkable for the LES-AV dataset, where RRWNet greatly outperforms the state-of-the-art methods in all metrics. The differences are, again, particularly pronounced for the topological metrics, with RRWNet achieving 47.05% COR and 51.68% INF for arteries (+35.31 pp and -36.14 pp over the second-best method, respectively) and 49.68% COR and 49.45% INF for veins (+23.77 pp and -24.46 pp over the second-best method, respectively). This showcases the generalization capabilities of RRWNet to different an unseen datasets, and its salient ability to generate topologically correct segmentation maps.
@@
Examples of segmentation maps obtained by the different models compared in Table 7 are shown in Figs. 9 and 10, for RITE and HRF datasets, respectively [^10] All the segmentation maps obtained by the models are publicly available at https://github.com/j-morano/rrwnet..
@@ 7 Conclusions
To rigorously assess the efficacy of the proposed framework, we implemented a straightforward instantiation based on the well-established U-Net architecture. This implementation was evaluated on the task of A/V segmentation and classification within several publicly available retinography image datasets. The quantitative results showed that the proposed method significantly outperformed the state of the art, both in terms of A/V classification accuracy and, more remarkably, in terms of topological consistency. The large improvement in topological consistency is particularly relevant because it opens up new applications in clinical practice, such as the precise computation of the AVR biomarker, where current state-of-the-art methods are deficient. In addition, the stand-alone application of the proposed RR module demonstrably improved the segmentation maps generated by all other methods compared, further substantiating its effectiveness as a generalizable post-processing step. To the best of our knowledge, this is the first robust post-processing method based solely on deep learning for A/V segmentation and classification.
@@
In conclusion, the proposed framework and its implementation represent an effective approach to A/V segmentation and classification, with the potential to be extended to other semantic segmentation tasks and modalities. We believe that this work will serve as a good reference implementation and benchmark, encourage further research in this direction, and contribute to the development of more robust and accurate semantic segmentation systems, with a particular focus on the field of ophthalmology.
@@ Acknowledgements
The financial support by the Austrian Federal Ministry for Labour and Economy and the National Foundation for Research, Technology and Development and the Christian Doppler Research Association is gratefully acknowledged.